# Un-Straightening Generative AI: How Queer Artists Surface and Challenge the Normativity of Generative AI Models

### v1

@@ 5.3. Un-Straightening Generative AI
As we described in Section 4.1.2, our participants critiqued the highly normative, stereotypical representations of queer people by GPT-4 and DALL-E 3 as thin, masculine, young, wealthy, Western, and ashamed of themselves. In some ways, our findings parallel prior work on biases embedded in generative AI models. For example, researchers have called attention to the gender and nationality biases in image models (Ghosh et al., 2024; Ghosh and Caliskan, 2023) and biases against queer couples in language models (Gillespie, 2024). However, other social biases our participants identified warrant greater research attention, such as those against fat people (Payne et al., 2023) and older adults. At the same time, our findings demonstrate that increased representation is not always desirable (DeVrio et al., 2022; Hoffmann, 2021), such as P15’s concerns about overly-diverse representations of police. As individuals’ diversity preferences may differ, designers could consider new user interaction paradigms (Morris, 2024) to better understand users’ preferences, such as asking clarifying questions before generating images of people.
@@
Although computing research on marginalized communities tends to focus on social biases and identity-based discrimination (Taylor et al., 2024), our participants’ concerns regarding GenAI extended beyond the representation of queer people in text and images. As we described above, the straight values explicitly and implicitly embedded in the design of GPT-4 and DALL-E 3 marginalized queer sexuality, politics, and style. In turn, we provided recommendations for GenAI developers to approach safety and style in ways that better align with our participants’ values. However, these implications for design and research still largely maintain the centralized power at the root of our participant’s critiques of GPT-4 and DALL-E 3, such as P13’s rhetorical question: ”Who defines what is Safe For Work?”
@@
Un-straightening GenAI requires shifting power away from the major corporations toward members of marginalized communities. Toward this, we encourage the participatory design of GenAI models outside corporate logics of scale (Young et al., 2024). At the same time, any truly meaningful participatory process must allow for the possibility that communities may simply not want to build GenAI models at all. In fact, numerous participants in our study were primarily interested in breaking — rather than making — GenAI models (Section 4.2.1). It follows that future research should support adversarial engagements between communities and those developing GenAI models, such as the development of tools to help artists protect their work from being used to train models without their consent (Shan et al., 2023, 2024). Whether contesting major corporate models or building alternatives, we caution those engaged in researcher-led initiatives from merely replacing corporate centralized power by centering themselves. Instead, we encourage researchers to engage in open-ended design (Sengers and Gaver, 2006), such as developing tools that individuals and communities can use to make/break GenAI models themselves (Li, 2024).
@@ 6. Conclusion
In this work, we examined the relationship between queer artists and GenAI through a medium-term workshop study. Our findings highlighted deep tensions between our participants’ values and the norms embedded in the design of GPT-4 and DALL-E 3. Despite these misalignment, our participants found queer ways to work around and with these straight models. While queer people have a long history of using technologies not designed with them in mind (Ahmed, 2019), our work highlights the limitations of dominant corporate models to meet the needs of queer artists. Instead, we call for a plurality GenAI models that embody community values throughout their design, development, and use.

### v2

@@ 5.3. Un-Straightening Generative AI
Our participants critiqued the highly normative, stereotypical representations of queer people by GPT-4 and DALL-E 3 as thin, masculine, young, wealthy, Western, and ashamed (Section 4.1.2). In some ways, our findings parallel prior work on biases embedded in GenAI models. Researchers have called attention to the gender and nationality biases in image models (Ghosh et al., 2024; Ghosh and Caliskan, 2023), biases against queer couples in language models (Gillespie, 2024), and moderation/representation biases against transgender non-cisgender identities in text-to-image models (Ungless et al., 2023). However, some biases our participants identified warrant greater research attention, such as those against fat people (Payne et al., 2023) and older adults. Our findings also demonstrate that increased representation is not always desirable (DeVrio et al., 2022; Hoffmann, 2021), such as P15’s concerns about overly-diverse police. As individuals’ diversity preferences may differ, designers could consider new user interaction paradigms (Morris, 2024) to better understand users’ preferences, such as asking clarifying questions before generating images of people.
@@
Although computing research on marginalized communities tends to focus on social biases and identity-based discrimination (Taylor et al., 2024b, a), our participants’ concerns regarding GenAI extended beyond the representation of queer people in text and images. As we described above, the straight values explicitly and implicitly embedded in the design of GPT-4 and DALL-E 3 marginalized queer sexuality, politics, and style. In turn, we provided recommendations for GenAI developers to approach safety and style in ways that better align with our participants’ values. However, these implications for design and research largely maintain the centralized power at the root of our participant’s critiques of GPT-4 and DALL-E 3, such as P13’s rhetorical question: ”Who defines what is Safe For Work?”
@@
Un-straightening GenAI requires shifting power away from major corporations toward marginalized communities (DeVrio et al., 2024; Sum and DeVrio, 2024). We encourage the participatory design of models outside corporate logics of scale (Young et al., 2024). At the same time, a meaningfully participatory process must allow for the possibility that communities may simply not want to build GenAI models at all. Numerous participants in our study were primarily interested in breaking — rather than making — GenAI models (Section 4.2.1). It follows that future research should support adversarial engagements between communities and those developing GenAI models, such as designing tools to help artists protect their work from being used to train models without their consent (Shan et al., 2023, 2024). Whether contesting major corporate models or building alternatives, we caution those engaged in researcher-led initiatives from merely replacing corporate centralized power by centering themselves. Instead, we encourage researchers to engage in open-ended design (Sengers and Gaver, 2006), such as making tools for individuals and communities to make/break GenAI models themselves (Li, 2024).
5.4. Limitations
Our work should not be interpreted as speaking for all ”queer artists.” Rather, our findings draw on the experiences of a particular group of queer artists. All our participants lived in the USA during the study and spoke English. This may limit the transferability of our findings, as conceptualizations of gender and sexuality are culturally situated (Nova et al., 2021; Grewal and Kaplan, 2001), GenAI models’ performance can differ between languages (Choudhury, 2023) and GenAI models tend to reproduce western biases (Mim et al., 2024). Future work should explore the experiences of queer artists outside the USA and in languages other than English.
Our work is also not intended to represent the diversity of queer artists within the USA. We tried to make our study more accessible by conducting workshops online and offering two weekly time slot options. However, those with limited internet access or less flexibility in their work schedule may have been unable to participate. This may be why, to the best of our knowledge, none of our participants lived in a rural areas. We caution against conflating queerness with urbanity (Hardy and Lindtner, 2017) and encourage future work on rural queer artists and technology. As conceptualizations of gender and sexuality are socially constructed (Taylor and Bruckman, 2024) and change over time (Foucault, 1990), our findings may not transfer to older queer communities. Our findings are also not only applicable to the experiences of queer artists: DALL-E 3’s gender moderation impacts those who are neither queer nor artists.
Some may take issue with the use of “queer” as non-normativity in queer theory (Ahmed, 2006). We do not suggest that queer people are all inherently non-normative. As noted by Grewal and Kaplan, ”queer subjects are not always already avant-garde for all time and in all places” (Grewal and Kaplan, 2001). Moreover, not every LGBTQ+ person necessarily identifies as queer. In this work, we specifically recruited participants who self-identified as “queer artists.” Our findings may have differed if we instead recruited LGBTQ+ identifying artists. Our work demonstrates the need to carefully distinguish between FAccT research using queer theory, about LGBTQ+ people, or (like our own) at the intersection of these discourses.
@@ 6. Conclusion
We examined the relationship between queer artists and GenAI through a medium-term workshop study, highlighting deep tensions between our participants’ values and the norms embedded in the design of GPT-4 and DALL-E 3. Nevertheless, our participants found queer ways to work around and with these straight models. While queer people have a long history of using technologies not designed with them in mind (Ahmed, 2019), our work highlights the limitations of dominant corporate models to meet the needs of queer artists. Instead, we call for a plurality GenAI models that embody community values throughout their design, development, and use.
7. Positionality
This research has been shaped by the positionalities of our research team. Multiple members of our research team identify as members of the LGBTQ+ community and all authors identify as white and/or East Asian. Our positionalities shape our analysis of queerness and queer identities through a largely western, academic lens. Our interpretations were also shaped by our academic disciplines as design, HCI, and NLP researchers as well as our experiences practicing and appreciating musical, visual, written, and textile arts. Lastly, our team is located at a university in a mid-size city in the United States of America. Our physical location shaped the geographic distribution of our participants and restricted our ability to recruit participants living outside the United States of America.
Acknowledgements.
First, we would like to thank our participants. We also thank Rose Chang, Daragh Byrne, William Agnew, Shivani Kapania, Cella Sum, Anh-Ton Tran, Ellen Simpson, Mary Gray and our reviewers for their helpful feedback. This work was supported, in part, by Microsoft Research’s Accelerating Foundation Models Research grant and Carnegie Mellon University’s Block Center for Technology and Society.